# Sparse Hyperparametric Itakura-Saito NMF via Bi-Level Optimization

### v2

@@ Abstract
The selection of penalty hyperparameters is a critical aspect in Nonnegative Matrix Factorization (NMF), since these values control the trade-off between the reconstruction accuracy and the adherence to desired constraints. In this work, we focus on an NMF problem involving the Itakura-Saito (IS) divergence, effective for extracting low spectral density components from spectrograms of mixed signals, enhanced with sparsity constraints. We propose a new algorithm called SHINBO, which introduces a bi-level optimization framework to automatically and adaptively tune the row-dependent penalty hyperparameters, enhancing the ability of IS-NMF to isolate sparse, periodic signals against noise. Experimental results showed SHINBO ensures precise spectral decomposition and demonstrates superior performance in both synthetic and real-world applications. For the latter, SHINBO is particularly useful, as noninvasive vibration-based fault detection in rolling bearings, where the desired signal components often reside in high-frequency subbands but are obscured by stronger, spectrally broader noise. By addressing the critical issue of hyperparameter selection, SHINBO advances the state-of-the-art in signal recovery for complex, noise-dominated environments.
@@
NMF, Itakura-Saito divergence, Hyperamater Optimization, Sparsity, Bi-level Optimization, Dynamical System
@@
[inst1]organization=Water Research Institute – National Research Council (IRSA-CNR), country=Italy
@@ 1 Introduction
Nonnegative Matrix Factorization (NMF) is a dimensionality reduction technique that approximates a nonnegative data matrix as the product of two (lower-dimensional) nonnegative matrices. A key challenge in NMF is setting penalty coefficients when additional constraints, such as sparsity or smoothness, are imposed [18]. These coefficients control the trade-off between reconstruction accuracy and constraint adherence, but their optimal values are highly dataset- and application-dependent, making the selection process non-trivial. For example, [6] proposes a variant of NMF that incorporates data-dependent penalties and introduces auxiliary constraints to enhance performance in tasks such as face recognition. Additionally, [20] presents multiplicative algorithms for NMF that enforce non-negativity and flow preservation constraints while introducing regularizations to ensure smoothness or sparsity. Finally, [11] adapts a minimization scheme for functions with non-differentiable constraints, known as PALM, to solve NMF problems, yielding solutions that can be both smooth and sparse—two highly desirable properties. In this work, we rely on prior studies [8, 9] in which the penalized problem is reformulated in a general form, and a strategy is proposed to tune the penalty coefficient automatically.
@@
with the objective function $D_{\beta}(\cdot,\cdot)$ being the $\beta$-divergence [13, 19], assessing the quality of the reconstruction $\bm{W}\bm{H}$ in fitting $\bm{X}$. We remind that the $\beta$-divergence for matrices is defined as $D_{\beta}(\bm{A},\bm{B})=\sum_{i}{\sum_{j}{d_{\beta}(a_{ij},b_{ij})}}$, where the function $d_{\beta}$ for each $x,y\in\mathbb{R}$ is defined as
$d_{\beta}(x,y)=\left\{\begin{array}[]{lc}\frac{1}{\beta(\beta-1)}(x^{\beta}+(\beta-1)y^{\beta}-\beta xy^{\beta-1})&\beta\in\mathbb{R}\setminus\{0,1\};\vspace{2mm}\\ x\log(\frac{x}{y})-x+y&\beta=1;\vspace{2mm}\\ \frac{x}{y}-\log(\frac{x}{y})-1&\beta=0.\end{array}\right.$
The function $\mathcal{P}:\mathbb{R}^{r\times n}\to\mathbb{R}$ is a penalty term on $\bm{H}$ that enforces a particular constraint. The vector $\bm{\lambda}\in\mathbb{R}^{r}_{+}$ in (1) contains a nonnegative penalty hyperparameters $[\lambda_{1},\dots,\lambda_{r}]$ associated to each row of $\bm{H}$.
In this study, we are interested in a special case of the $\beta$-divergence. When $\beta=0$, the $\beta$-divergence boils down to the Itakura-Saito (IS) divergence which is a scale-invariant measure of dissimilarity, useful for extracting low spectral density components from spectrograms of mixed signals. In many practical applications, the desired components are located in a high frequency subband and are often masked by much stronger and spectrally wider noisy disturbances. This is the case of a noninvasive vibration-based fault detection in rolling bearings [23, 24, 25]. The vibrational diagnostic signals measured from faulty bearings on the laboratory test rig [16] are used in our study to validate the proposed algorithm.
The signal of interest (SOI) in this application is represented by a periodic and impulsive signal. The observed signal should contain the SOI and other perturbations, usually expressed by a strong independent and identically distributed (i.i.d.) Gaussian noise. Applying NMF to the spectrogram of the measured signal, we expect that the representative of the SOI will be expressed by one of the temporal components of NMF, i.e., one row of matrix $\bm{H}$. This component, which presents a periodic spiky signal, is obviously very sparse. The other components, which represent the noisy perturbations, should not be sparse. To enforce NMF to search for the desired component, we introduced the penalty for the rows of $\bm{H}$ using the term $\mathcal{P}(\textrm{Diag}(\bm{\lambda})\bm{H})$ in (1). The proposed bi-level approach should perform data-driven adaptation of the hyperparameters (vector $\bm{\lambda}$) to the desired nature of the estimated components.
@@
The contribution of this work is two folded.
@@
    • (a)
      The penalty hyperparameter is not known in advance. This parameter is formulated as an optimization variable (in form of bi-level optimization model) and is solved by a bi-level optimization method.
    • (b)
      The penalty hyperparameter is row-dependent. Note that Problem (1) is not the standard penalized NMF [18], which applies the same penalty coefficient to all rows in the matrix $\bm{H}$. In Problem (1) each row of $\bm{H}$ is penalized by its own penalty parameter.
@@
    New algorithm. For Problem (1), we present a new multiplicative update (see Equation (12)), and a way to automatically tune the penalty hyperparameter based on bi-level strategy (see details in Section 3).
@@
The symbol $\bm{0}$ denote the zero matrix, the symbol $\bm{E}_{a\times b}$ is all-one matrix sizing $a\times b$ with $a,b\in\mathbb{N}$. The notation $\bm{v}$ denotes a column vector and the notation $\underline{\bm{v}}$ means $\bm{v}$ is a row vector.
On matrix, $\bm{A}^{\top}$ is the transpose of $\bm{A}$, and $\bm{A}^{2}=\bm{A}\bm{A}$. The symbol $\bm{A}\odot\bm{B}$ refers to the Hadamard (element-wise) product between $\bm{A}$ and $\bm{B}$ of conformal dimensions, and the symbol $\bm{A}\oslash\bm{B}$ with $\bm{B}\neq\bm{0}$ refers to the Hadamard division, and we denote $\bm{A}^{\odot k}$ as the Hadamard power-$k$ of $\bm{A}$.
Given $n\in\mathbb{N}$, we denote $[n]\coloneqq\{1,2,\ldots,n\}$. The symbols $k,t\in\mathbb{N}$ indicate iteration counters. The symbol $\bm{A}^{k}$ refers to the variable $\bm{A}$ at the iteration $k$, $A_{ij}$ or $a_{ij}$ is the $(i,j)$-th element of $\bm{A}$. Lastly $\bm{A}_{i:}$ and $\bm{A}_{:j}$ are the $i$-th and $j$-th row and column of $\bm{A}$, respectively.
[…]
In this section, we discuss the main focus of the paper, Problem (1) and the overall framework of the proposed algorithm.
@@
We focus on Problem (1) with $\beta=0$. As $\mathcal{P}$ we chose a particular penalty function, effective in increasing sparsity, that is the diversity measure $J$ [7, 9]. In a particular case, if the matrix is nonnegative, the diversity measure can be written in terms of the trace operator as
@@
where $\mathbf{E}$ is the squared matrix of ones. Thus, by properties of trace, Problem (1) becomes
@@
[…]
1 Input: $\bm{X}\in\mathbb{R}^{m\times n}_{+}$ and factorization rank $r$.
@@
4       $\bm{W}^{k}=\text{update}(\bm{X},\bm{W}^{k-1},\bm{H}^{k-1})$ classic MU-update
@@
6             $\underline{\bm{h}}_{l}^{k-1,0}=\underline{\bm{h}}_{l}^{k-1},\lambda^{k-1,0}_{l}=\lambda^{k-1}_{l}$ initialization
@@
9                   $\dfrac{\partial\mathcal{R}}{\partial\lambda_{l}}$ as in (10) % hypergradient
@@
11            $\bm{\lambda}^{k,T}=\text{update}(\bm{\lambda}^{k,T},\nabla_{\bm{\lambda}}\mathcal{R}(\bm{\lambda}))$ projected gradient update
@@
[…]
@@
The next section introduces the update on $\bm{H}$ and $\bm{\lambda}$ by a bi-level method, which is the main contribution of the paper.
@@ 3 Bi-level Optimization for the subproblem
Section organization and general overview of the approach. Under a fixed and given $\bm{W}^{k}$, the goal is to obtain an updated version of $\bm{H}$ and $\bm{\lambda}$ that approximately solves Problem (4). The bi-level approach has the following steps:
@@
    First we replace the constrained optimization problem (4) by a bi-level optimization problem for $(\underline{\bm{h}}_{l},\lambda_{l})$, with $l\in[r]$, see (5).
@@
    Lastly we use the hypergradient to obtain the solution $\bm{\lambda}$ by a gradient descent approach. See (6).
@@ 1. Bi-level formulation
In Algorithm 1, the update of $\bm{H}$ and $\bm{\lambda}$ is performed in $r$ steps, where each step is aimed to update the $l$-th component $(\underline{\bm{h}}_{l},\lambda_{l})$. This is achieved by solving the following bi-level problem
$\begin{array}[]{rcl}\displaystyle\min_{\lambda_{l}\geq 0}\left\{\begin{array}[]{rl}\displaystyle\mathcal{r}(\lambda_{l})=&\displaystyle\inf_{\underline{\bm{h}}_{l}(\lambda_{l})}D_{2}\big{(}\bm{X},\bm{R}+\bm{w}_{l}\underline{\bm{h}}_{l}(\lambda_{l})\big{)}\qquad\text{(OP)}\vspace{2mm}\\ &\,\textrm{s.t.}\,\,\,\displaystyle\underline{\bm{h}}_{l}(\lambda_{l})\in\operatorname*{\textrm{argmin}\,}_{\bm{u}\in\mathbb{R}^{n}_{+}}D_{0}(\bm{X},\bm{R}+\bm{w}_{l}\underline{\bm{u}})+\lambda_{l}^{2}\|\underline{\bm{u}}\|_{1}^{2}\quad\text{(IP)}\end{array}\right\},\end{array}$ (5)
where the matrix $\bm{R}$ is the residual obtained isolating in $\bm{W}\bm{H}$ the $l$-th component of $\bm{w}_{l}\in\mathbb{R}^{m}$ (column of $\bm{W}$) and $\underline{\bm{h}}_{l}\in\mathbb{R}^{n}$ (row of $\bm{H}$), which is
@@
The function $\mathcal{r}:\mathbb{R}\to\mathbb{R}$ is called Response function of the outer problem related to $\underline{\bm{h}}_{l}$. In the outer problem, the objective $D_{2}$ is the $\beta$-divergence with $\beta=2$, which is the Frobenius norm. The inner problem is represented by the $\beta$-divergence with $\beta=0$, which is the Itakura-Saito divergence $D_{0}$, regularized by the row-wise diversity measure defined in (2).
@@
Note that the squared $\ell_{1}$-norm in the inner problem on $\bm{u}$ can be seen as a non-smooth regularization term and therefore proximal gradient descent can be applied on $\bm{u}$ (see [10] and [1, Lemma 6.70]), however, their approach is applied to convex objective function, where here in (5) the objective function (the IS-divergence) is possibly nonconvex [12] thus proximal gradient descent do not have convergence guarantee.

### v3

# Sparse Hyperparametric Itakura-Saito Nonnegative Matrix Factorization via Bi-Level Optimization

@@ Abstract
The selection of penalty hyperparameters is a critical aspect in Nonnegative Matrix Factorization (NMF), since these values control the trade-off between reconstruction accuracy and adherence to desired constraints. In this work, we focus on an NMF problem involving the Itakura-Saito (IS) divergence, which is particularly effective for extracting low spectral density components from spectrograms of mixed signals, and benefits from the introduction of sparsity constraints. We propose a new algorithm called SHINBO, which introduces a bi-level optimization framework to automatically and adaptively tune the row-dependent penalty hyperparameters, enhancing the ability of IS-NMF to isolate sparse, periodic signals in noisy environments. Experimental results demonstrate that SHINBO achieves accurate spectral decompositions and demonstrates superior performance in both synthetic and real-world applications. In the latter case, SHINBO is particularly useful for noninvasive vibration-based fault detection in rolling bearings, where the desired signal components often reside in high-frequency subbands but are obscured by stronger, spectrally broader noise. By addressing the critical issue of hyperparameter selection, SHINBO improves the state-of-the-art in signal recovery for complex, noise-dominated environments.
@@
Nonnegative Matrix Factorization, Itakura-Saito divergence, Hyperameter Optimization, Sparsity, Bi-level Optimization, Dynamical System
@@
[inst1]organization=Water Research Institute - National Research Council (IRSA-CNR), country=Italy
@@ 1 Introduction
Nonnegative Matrix Factorization (NMF) is a dimensionality reduction technique that approximates a nonnegative data matrix as the product of two (lower-dimensional) nonnegative matrices. A key challenge in NMF is to determine penalty coefficients when additional constraints, such as sparsity or smoothness, are imposed [1]. These coefficients control the trade-off between reconstruction accuracy and constraint adherence, but their optimal values are highly dependent on both the dataset and the specifict application, making the selection process non-trivial. For example, [2] proposes a variant of NMF that incorporates data-dependent penalties and introduces auxiliary constraints to enhance performance in tasks such as face recognition. Additionally, [3] presents multiplicative algorithms for NMF that enforce non-negativity and flow preservation constraints while introducing regularizations to ensure smoothness or sparsity. Finally, [4] adapts a minimization scheme for functions with non-differentiable constraints, known as PALM, to solve NMF problems, yielding solutions that can be both smooth and sparse—two highly desirable properties. In this work, we rely on previous studies [5, 6] in which the penalized problem is reformulated in a general form, and a strategy is proposed to automatically adjust the penalty coefficient.
@@
with the objective function $D_{\beta}(\cdot,\cdot)$ being the $\beta$-divergence [7, 8], assessing the quality of the reconstruction $\bm{W}\bm{H}$ in fitting $\bm{X}$  [^1] We remind that the $\beta$-divergence for matrices can be computed element-wise as $D_{\beta}(\bm{A},\bm{B})=\sum_{i}{\sum_{j}{d_{\beta}(a_{ij},b_{ij})}}$, where the function $d_{\beta}$ for each $x,y\in\mathbb{R}$ is defined as $d_{\beta}(x,y)=\left\{\begin{array}[]{lc}\frac{1}{\beta(\beta-1)}(x^{\beta}+(\beta-1)y^{\beta}-\beta xy^{\beta-1})&\beta\in\mathbb{R}\setminus\{0,1\};\vskip 5.69054pt\\ x\log(\frac{x}{y})-x+y&\beta=1;\vskip 5.69054pt\\ \frac{x}{y}-\log(\frac{x}{y})-1&\beta=0.\end{array}\right$. . The function $\mathcal{P}:\mathbb{R}^{r\times n}\to\mathbb{R}$ is a penalty term on $\bm{H}$ that enforces a particular constraint. The vector $\bm{\lambda}\in\mathbb{R}^{r}_{+}$ in (1) contains nonnegative penalty hyperparameters $\lambda_{1},\dots,\lambda_{r}]$ associated to each row of $\bm{H}$.
In this study, we focus on a special case of the $\beta$-divergence within NMF. The $\beta$-divergence offers a tunable objective function that adapts to the statistics of the data: by varying $\beta$ it interpolates between the Euclidean distance ($\beta=2$), the Kullback-Leibler divergence ($\beta=1$), and—in the limit $\beta\to 0$—the Itakura-Saito (IS) divergence. The IS divergence is scale-invariant and normalizes the reconstruction error by the signal magnitude, so weak spectral components (i.e. frequency components with low spectral density) contribute proportionally to the objective rather than being dominated by high-energy regions. As a results, IS-NMF is well suited to preserving fine, low-energy structures (e.g., weak higher-order harmonics, background noise, or faint events) that are often missed by Euclidean/KL objectives. Moreover, IS connects naturally to a multiplicative Gamma-noise model for power spectra, which aligns with the heavy-tailed, non-Gaussian behavior commonly encountered in audio and vibration spectrograms.
Our target application is non-invasive, vibration-based fault detection in rolling bearings [9, 10, 11, 12, 13]. The desired signal of interest (SOI) is periodic and impulsive, yet it typically resides in a high-frequency subband and is masked by much stronger, spectrally broad disturbances. We validate our approach on laboratory test-rig measurements of faulty bearings [14]. Although time-domain perturbations are often modeled as additive i.i.d. Gaussian noise, after time-frequency transformation, the IS divergence provides a statistically well-matched criterion for factorizing spectrograms with multiplicative fluctuations.
Related work has applied $\beta$-NMF to source separation in audio [15] and established identifiability via minimum-volume (MinVol) regularization. The “sufficiently scattered” condition [16] relaxes separability assumptions, and subsequent studies showed that minimizing the convex-hull volume of basis columns can recover factors even in rank-deficient or nonseparable regimes [15], thereby linking MinVol-NMF to practical uniqueness guarantees.
Methodologically, we apply NMF to the spectrogram of the measured signal. Let $\bm{H}$ denote the temporal activations. We expect the SOI to be captured by one row of $\bm{H}$ that exhibits sparse, periodic, spike-like activity, whereas rows associated with broadband disturbances are comparatively dense. To bias the factorization toward the SOI, we impose a row-sparsity-promoting penalty on $\bm{H}$ via the term $\mathcal{P}(\textrm{Diag}(\bm{\lambda})\bm{H})$ in (1). We further adopt a bi-level scheme that adapts the hyperparameters $\bm{\lambda}$ from the data, aligning the penalty strength with the latent structure of the estimated components.
@@
The contribution of this work is twofold.
@@
    • The penalty hyperparameter is not known in advance. We treat this parameter as an optimization variable, embedding it in a bi-level optimization framework and solving it by a bi-level optimization method.
    • The penalty hyperparameter is row-dependent. Note that Problem (1) is different from standard penalized NMF [1], which applies the same penalty coefficient to all rows in the matrix $\bm{H}$. In Problem (1), each row of $\bm{H}$ is penalized by its own penalty parameter.
@@
    New algorithm. For Problem (1), we present a new multiplicative update (see Equation (12)), and a way to automatically tune the penalty hyperparameter based on a bi-level strategy (see details in Section 3).
@@
The symbol $\bm{0}$ denotes the zero matrix, the symbol $\bm{E}_{a\times b}$ is the all-in matrix size $a\times b$ with $a,b\in\mathbb{N}$. The notation $\bm{v}$ denotes a column vector and the notation $\underline{\bm{v}}$ means that $\bm{v}$ is a row vector. On matrix, $\bm{A}^{\top}$ is the transpose of $\bm{A}$, and $\bm{A}^{2}=\bm{A}\bm{A}$. The symbol $\bm{A}\odot\bm{B}$ refers to the Hadamard (element-wise) product between $\bm{A}$ and $\bm{B}$ of conformal dimensions, and the symbol $\bm{A}\oslash\bm{B}$ with $\bm{B}\neq\bm{0}$ refers to the Hadamard division, and we denote $\bm{A}^{\odot k}$ as the Hadamard power-$k$ of $\bm{A}$. Given $n\in\mathbb{N}$, we denote $n]\coloneqq\{1,2,\ldots,n\}$. The symbols $k,t\in\mathbb{N}$ indicate iteration counters. The symbol $\bm{A}^{k}$ refers to the variable $\bm{A}$ at the iteration $k$, $A_{ij}$ or $a_{ij}$ is the $(i,j)$th element of $\bm{A}$. Lastly $\bm{A}_{i:}$ and $\bm{A}_{:j}$ are the $i$th and $j$th row and column of $\bm{A}$, respectively.
2 The overall optimization framework of $\operatorname*{\texttt{SHINBO}}$
In this section, we discuss the focus of the paper, Problem (1) and the overall framework of the proposed algorithm.
@@ 2 The overall optimization framework of $\operatorname*{\texttt{SHINBO}}$
We focus on Problem (1) with $\beta=0$. As $\mathcal{P}$ we chose a particular penalty function, effective in increasing sparsity, that is the diversity measure $J$ [17, 6]. In a particular case, if the matrix is nonnegative, the diversity measure can be written as
@@
where $\mathbf{E}$ is the matrix of 1s. By properties of trace, Problem (1) becomes
@@
The proposed optimization algorithm $\operatorname*{\texttt{SHINBO}}$
We propose an algorithm, called $\operatorname*{\texttt{SHINBO}}$, to find a local minima for Problem (3) as follows.
1Input: $\bm{X}\in\mathbb{R}^{m\times n}_{+}$ and factorization rank $r$.
@@
4 $\bm{W}^{k}=\text{update}(\bm{X},\bm{W}^{k-1},\bm{H}^{k-1})$ % classic MU-update
@@
6    $\underline{\bm{h}}_{l}^{k-1,0}=\underline{\bm{h}}_{l}^{k-1},\lambda^{k-1,0}_{l}=\lambda^{k-1}_{l}$ % initialization
@@
9       $\frac{\partial\mathcal{R}}{\partial\lambda_{l}}$ as in (10) % hypergradient
@@
11    $\bm{\lambda}^{k,T}=\text{update}(\bm{\lambda}^{k,T},\nabla_{\bm{\lambda}}\mathcal{R}(\bm{\lambda}))$ % projected gradient update
@@
Algorithm 1 $\operatorname*{\texttt{SHINBO}}$
Note that $\operatorname*{\texttt{SHINBO}}$ is composed by two main parts: one devoted to the update of $\bm{W}$ (reviewed in the following) and the other one based on bi-level strategy to optimize simultaneously on $\bm{H}$ and $\bm{\lambda}$.
@@
The following section introduces the bi-level update on $\bm{H}$ and $\bm{\lambda}$ by a bi-level method, which is the main contribution of the paper.
@@ 3 Bi-level Optimization for the subproblem
Section organization and overview of the approach. Under a fixed and given $\bm{W}^{k}$, the goal is to obtain an updated version of $\bm{H}$ and $\bm{\lambda}$ that approximately solves Problem (4). The bi-level approach has the following steps:
@@
    We first rewrite the constrained optimization problem (4) by a bi-level optimization problem for $(\underline{\bm{h}}_{l},\lambda_{l})$, with $l\in[r]$, see (5).
@@
    Lastly, we use the hypergradient to obtain the solution $\bm{\lambda}$ by a gradient descent approach. See (6).
@@ 1. Bi-level formulation
In Algorithm 1, the update of $\bm{H}$ and $\bm{\lambda}$ is performed in $r$ steps, where each step is aimed to update the $l$th component $(\underline{\bm{h}}_{l},\lambda_{l})$. This is achieved by solving the following bi-level problem
$\begin{array}[]{rcl}\hskip-14.22636pt\displaystyle\min_{\lambda_{l}\geq 0}\left\{\hskip-4.2679pt\begin{array}[]{rl}\displaystyle\mathcal{r}(\lambda_{l})=&\hskip-8.53581pt\displaystyle\inf_{\underline{\bm{h}}_{l}(\lambda_{l})}D_{2}\big(\bm{X},\bm{R}+\bm{w}_{l}\underline{\bm{h}}_{l}(\lambda_{l})\big)\qquad\text{(OP)}\vskip 2.84526pt\\ &\hskip-8.53581pt\,\textrm{s.t.}\,\,\,\displaystyle\underline{\bm{h}}_{l}(\lambda_{l})\in\operatorname*{\textrm{argmin}\,}_{\bm{u}\in\mathbb{R}^{n}_{+}}D_{0}(\bm{X},\bm{R}+\bm{w}_{l}\underline{\bm{u}})+\lambda_{l}^{2}\|\underline{\bm{u}}\|_{1}^{2}\quad\text{(IP)}\end{array}\hskip-5.69054pt\right\},\end{array}$ (5)
where the matrix $\bm{R}$ is the residual obtained isolating in $\bm{W}\bm{H}$ the $l$th component of $\bm{w}_{l}\in\mathbb{R}^{m}$ (column of $\bm{W}$) and $\underline{\bm{h}}_{l}\in\mathbb{R}^{n}$ (row of $\bm{H}$), which is
@@
The function $\mathcal{r}:\mathbb{R}\to\mathbb{R}$ is called the Response function of the outer problem related to $\underline{\bm{h}}_{l}$. In the outer problem, the objective $D_{2}$ is the $\beta$-divergence with $\beta=2$, which is the Frobenius norm. The inner problem is represented by the $\beta$-divergence with $\beta=0$, which is the Itakura-Saito divergence $D_{0}$, regularized by the row-wise diversity measure defined in (2).
@@
Note that the squared $\ell_{1}$-norm in the inner problem on $\bm{u}$ can be seen as a non-smooth regularization term and therefore proximal gradient descent can be applied on $\bm{u}$ (see [19] and [20, Lemma 6.70]), however, their approach is applied to convex objective function, where here in (5) the objective function (the IS-divergence) is possibly nonconvex [18] thus proximal gradient descent do not have convergence guarantee.